OpenStudy (anonymous):

(sinx/1-cosx)+(sinx/1+cosx)=2cscx

OpenStudy (kc_kennylau):

\[\frac{\sin x}{1-\cos x}+\frac{\sin x}{1+\cos x}=2\csc x\]

OpenStudy (kc_kennylau):

Finally a man here who knows how to use parentheses

OpenStudy (anonymous):

what exactally are you looking for?

OpenStudy (anonymous):

Verify the identity

OpenStudy (kc_kennylau):

lol my fault

OpenStudy (kc_kennylau):

Okay then, combine the fractions together first

OpenStudy (anonymous):

2 sin(x) = 2 csc(x)

OpenStudy (anonymous):

I know it's combing denominators first; it is just after that I need.

OpenStudy (kc_kennylau):

@OG_UnLeAsH: You forgot the denominator

OpenStudy (anonymous):

darn it

OpenStudy (kc_kennylau):

So this identity is correct

OpenStudy (kc_kennylau):

okay @ayolisse can you draw out what you got after combining?

OpenStudy (anonymous):

posted on wrong question whoops

OpenStudy (anonymous):

sinx(1+cosx)+sinx(1-cosx)/(1-cosx)(1+cosx)

OpenStudy (kc_kennylau):

Distribute?

OpenStudy (anonymous):

\[sinx+sinxcosx+sinx-sinxcosx/ (1-cosx)(1+cosx)\]

OpenStudy (kc_kennylau):

Add the like terms

OpenStudy (anonymous):

\[2sinx+1-cosxsinx/ (1-cosx)(1+cosx)\]

OpenStudy (kc_kennylau):

Where did the 1 come from?

OpenStudy (anonymous):

\[(\sin^2)x+(\cos^2)x=1\]

OpenStudy (kc_kennylau):

But there's no sin^2x or cos^2x

OpenStudy (anonymous):

then what would you do cause now I am lost

OpenStudy (kc_kennylau):

Continue with \(\dfrac{\sin x+\sin x\cos x+\sin x-\sin x\cos x}{(1-\cos x)(1+\cos x)}\)

OpenStudy (anonymous):

\[2sinx+cosxsinx-cosxsinx/(1-cosx)(1+cosx)\]

OpenStudy (kc_kennylau):

And then?

OpenStudy (anonymous):

do the cosxsinx cancel?

OpenStudy (kc_kennylau):

Yes

OpenStudy (anonymous):

now what about the denominator @kc_kennylau

OpenStudy (kc_kennylau):

Distribute

OpenStudy (anonymous):

\[2sinx/1+cosx-cosx-(\cos^2)x\]

OpenStudy (anonymous):

the cosx cancel right?

OpenStudy (kc_kennylau):

Yep

OpenStudy (anonymous):

then what do you do after that?

OpenStudy (kc_kennylau):

Do you remember the identity you've said in the wrong time?

OpenStudy (anonymous):

the (sin^2)x+(cos^2)x=1 ?

OpenStudy (kc_kennylau):

Yep

OpenStudy (kc_kennylau):

Make cos^2x the subject

OpenStudy (anonymous):

do you make the denominator sin^2x?

OpenStudy (kc_kennylau):

Yes

OpenStudy (anonymous):

okay i go the answer